# Orthogonal polynomials for the singularly perturbed Laguerre weight, Hankel determinants and asymptotics

### v1

@@ Abstract
Based on the work of Chen and Its [J. Approx. Theory 162 (2010) 270–297], we further study orthogonal polynomials with respect to the singularly perturbed Laguerre weight $w(x;t,\alpha)={x^{\alpha}}{\mathrm{e}^{-x-\frac{t}{x}}},\;x\in\mathbb{R}^{+},\;\alpha>-1,\;t\geq 0$. By using the ladder operators and associated compatibility conditions for orthogonal polynomials with general Laguerre-type weights, we derive the second-order differential equation satisfied by the orthogonal polynomials, a system of difference equations and a system of differential-difference equations for the recurrence coefficients. We also investigate the properties of the zeros of the orthogonal polynomials. Using Dyson’s Coulomb fluid approach together with the discrete system, we obtain the large $n$ asymptotic expansions of the recurrence coefficients, the sub-leading coefficient of the monic orthogonal polynomials, the Hankel determinant and the normalized constant for fixed $t>0$. It is found that all the asymptotic expansions are singular at $t=0$. We also study the long-time ($t\rightarrow\infty$) asymptotics of these quantities explicitly for fixed $n\in\mathbb{N}$ from the Toda-type system.
@@ 1 Introduction
In the important work [12], Chen and Its studied a singular linear statistic of the Laguerre unitary ensemble (LUE) inspired by an integrable quantum field theory at finite temperature [28]. This problem is equivalent to the study of the Hankel determinant and orthogonal polynomials for a singularly perturbed Laguerre weight
@@
The matrix model and Hankel determinant associated with the weight (1.1) have played an important role in the study of statistics for zeros of the Riemann zeta function [6] and the bosonic replica field theories [37]. Orthogonal polynomials and Hankel determinants for singularly perturbed Gaussian and Jacobi weights have also been extensively studied in the past few years; see, e.g., [7, 9, 31, 35]. Matrix models with higher-order poles have been investigated in [1, 17, 18].
We would like to mention that orthogonal polynomials and Hankel determinants for the deformed measure $\mathrm{e}^{\frac{t}{x}}d\mu(x)$ have a very close relationship with the peakon solutions of the Camassa-Holm equation [3, 4, 5]. Orthogonal polynomials and Hankel determinants for the singularly perturbed Laguerre weight (1.1) have been used to study the long-time asymptotics of the peakon solutions of the Camassa-Holm equation recently by Chang, Eckhardt and Kostenko [8, Section 6].
@@
where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma\leq 2$ and deg $\tau$=1. For semi-classical orthogonal polynomials, they have a weight $w(x)$ that satisfies the Pearson equation (1.2), where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma>2$ or deg $\tau\neq 1$. See, e.g., [41, Section 1.1.1]. It is not difficult to find that (1.1) is a semi-classical weight, since it satisfies the Pearson equation (1.2) with
@@
Please note that we use the notation $\alpha_{n},\beta_{n}$ for the recurrence coefficients (since this is standard), which should not cause confusion with the Laguerre weight parameter $\alpha$ as the latter is never subscripted with $n$. From (1.3), (1.4) and (1.5), the recurrence coefficients also have the following expressions:
@@
where $K_{\nu}(z)$ is the modified Bessel function of the second kind of order $\nu$ and it has the integral representation [36, Equation (10.32.10)]
@@
The rest of the paper is organized as follows. In Section 2, by applying the ladder operators and compatibility conditions for the generalized Laguerre-type weight to our problem, we show that the orthogonal polynomials satisfy a second-order differential equation, and derive a discrete system and a Toda-type system for the recurrence coefficients. We also present the relations among the logarithmic derivative of the Hankel determinant, the sub-leading coefficient of the monic orthogonal polynomials, the recurrence coefficients and the auxiliary quantities. In Section 3, we obtain a mixed recurrence relation for the orthogonal polynomials and then study some properties of the zeros of the orthogonal polynomials. In Section 4, we investigate the large $n$ asymptotics of the recurrence coefficients $\alpha_{n}(t,\alpha)$ and $\beta_{n}(t,\alpha)$, the sub-leading coefficient $\mathrm{p}(n,t,\alpha)$, the Hankel determinant ${D_{n}}(t,\alpha)$ and the normalized constant ${h_{n}}(t,\alpha)$} for fixed $t>0$. We discuss the long-time ($t\rightarrow\infty$) asymptotics of these quantities explicitly for fixed $n\in\mathbb{N}$ in Section 5.
@@ 2 Ladder operators, difference and differential equations
where $w_{0}(x)$ is continuously differentiable on $0,\infty)$ and all the moments of $w(x)$ exist. (Chen and Ismail’s ladder operator method requires $\alpha>0$.) For our problem, the monic orthogonal polynomials $P_{n}(x)$ satisfy the ladder operator equations
@@
$B_{n}^{2}(x)+\mathrm{v}^{\prime}(x)B_{n}(x)+\sum_{j=0}^{n-1}A_{j}(x)=\beta_{n}A_{n}(x)A_{n-1}(x).$ ($S_{2}^{\prime}$)
@@
The expressions of ${A_{n}}(x)$ and ${B_{n}}(x)$ in the above lemma are coincident with [12, Lemma 2]. It can be seen that our derivation is more straightforward and the results are valid for $\alpha>-1,\;t\geq 0$.
@@
The auxiliary quantity $R_{n}(t)$, related to the recurrence coefficient $\alpha_{n}(t)$ via (2.17), satisfies the Painlevé III  equation [29, (A.45.4 )]
@@
Then, the equation (2.35) is converted into the $\sigma$-form of the Painlevé III equation [26, (C.29)]
@@ 3 Zeros of orthogonal polynomials with the singularly perturbed Laguerre weight
$+\left(1-\frac{e_{n}}{\beta_{n-1}(t,\alpha)}\right)P_{n}(x;t,\alpha),\qquad n=2,3,\ldots,$ (3.1)
@@
Following [16, Lemma 3.6], we have for our weight
@@
By using Christoffel’s formula (see, e.g., [24, Theorem 2.7.1]) and letting
@@
Let $x_{1,n}<x_{2,n}<\cdots<x_{n,n}$ denote the $n$ zeros of the orthogonal polynomials $P_{n}(x;t,\alpha)$ with respect to the weight (1.1). Then the zeros
@@
    $0<x_{1,n}<\alpha_{n-1}+\frac{d_{n}\beta_{n-1}}{e_{n}}<x_{n,n},\qquad n=2,3,\ldots,$ (3.5)
@@
    are both increasing functions of $x$ on $\mathbb{R}^{+}$. Using Markov’s monotonicity theorem [39, Theorem 6.12.1], we have that the $j$th zero $x_{j,n}$ (for fixed $j=1,2,\ldots,n$) is an increasing function of both $t$ and $\alpha$.
@@
    which corresponds to $k=1$ and $G_{1}(x)=x-\alpha_{n-1}(t,\alpha)$ in [21, Equation (1)]. It follows from [21, Corollary 2.2] that $\alpha_{n-1}(t,\alpha)$ is an inner bound of the extreme zeros $x_{1,n}$ and $x_{n,n}$. The outer bounds $a_{n}$ and $b_{n}$ for the extreme zeros come from Ismail [24, Theorems 7.2.6 and 7.2.7] (see also Ismail and Li [25, Theorems 2 and 3]) using chain sequences.
@@
    Using the mixed recurrence relation (3.1) and following the similar analysis in [16, Theorem 3.9], it can be found that the zero of $G(x):=\frac{e_{n}}{\beta_{n-1}}\left(x-\alpha_{n-1}\right)-d_{n}$ (the coefficient of $P_{n-1}(x)$ in (3.1)) and the n2n-2 zeros of $P_{n-2}(x;t,\alpha+2)$ interlace with the $n$ zeros of $P_{n}(x;t,\alpha)$. It follows that the zero $\alpha_{n-1}+\frac{d_{n}\beta_{n-1}}{e_{n}}$ of $G(x)$ must lie between the two extreme zeros of $P_{n}(x;t,\alpha)$. Hence, we obtain another inner bound for the extreme zeros in (3.5).∎
@@
If $t=0$, our orthogonal polynomials $P_{n}(x;t,\alpha)$ are reduced to the classical (monic) Laguerre polynomials and $\alpha_{n}(0,\alpha)=2n+1+\alpha,\;\beta_{n}(0,\alpha)=n(n+\alpha)$. In this case, Theorem 3.2(iii) coincides with the result for classical Laguerre polynomials [24, Theorem 7.2.8] by choosing $c_{n}=4$ for $a_{n}$ and $c_{n}=4\cos^{2}\left(\frac{\pi}{n+1}\right)$ for $b_{n}$.
@@ 4 Large $n$ asymptotics
In the framework of Dyson’s Coulomb fluid approach [22], the eigenvalues of the Hermitian matrices from the unitary ensemble can be approximated by a continuous fluid with an equilibrium density when $n$ is sufficiently large. For our problem, it can be seen that the potential $\mathrm{v}(x)$ in (2.6) satisfies the condition that $x\mathrm{v}^{\prime}(x)$ is increasing on $\mathbb{R}^{+}$. This leads to the fact that the support of the equilibrium density is a single interval, denoted by (a,b),a>0(a,b),\;a>0; see [38, p. 199].
@@
Following the similar computations in [32, p. 406–407], multiplying by $\frac{1}{\sqrt{(b-x)(x-a)}}$ on both sides of (4.3) and integrating with respect to $x$ from $a$ to $b$, we have
@@
Our results for the large $n$ asymptotic expansions of the recurrence coefficients and the quantity $H_{n}(t)$ in Theorems 4.1 and 4.3 coincide with those in the work of Xu, Dai and Zhao [43, Corollary 1] using the Deift-Zhou steepest descent method for Riemann-Hilbert problems. Our method has the advantage of finding as many terms as one wants for the expansions. Furthermore, it can be seen that the large $n$ asymptotic expansions of the recurrence coefficients and the quantity $H_{n}(t)=t\frac{d}{{dt}}\ln{D_{n}}(t)$ are singular at the origin, which implies that the large $n$ asymptotic expansion of the Hankel determinant $D_{n}(t)$ can not be derived directly by integrating the large $n$ asymptotic expansion of $H_{n}(t)/t$ from 0 to $t$. This shows that Theorem $1$ in [43] is not correct. We derive the asymptotic expansion of the Hankel determinant in the following theorem.
@@ 5 Long-time asymptotics
The long-time asymptotics (without higher-order terms) of the Hankel determinant in (5.7) has also been recently derived by Chang, Eckhardt and Kostenko [8, Lemma 6.7] employing different method, which has been used to obtain the long-time asymptotics of the peakon solutions of the Camassa-Holm equation. In addition, the constant (5.8) also appeared in the asymptotic expansion of the Hankel determinant for the generalized Airy weight as $t\to+\infty$ [34].

### v2

@@ Abstract
Based on the work of Chen and Its [J. Approx. Theory 162 (2010) 270–297], we further study orthogonal polynomials with respect to the singularly perturbed Laguerre weight $w(x;t,\alpha)={x^{\alpha}}{\mathrm{e}^{-x-\frac{t}{x}}},\;x\in\mathbb{R}^{+},\;\alpha>-1,\;t\geq 0$. By using the ladder operators and associated compatibility conditions for orthogonal polynomials with general Laguerre-type weights, we derive the second-order differential equation satisfied by the orthogonal polynomials, a system of difference equations and a system of differential-difference equations for the recurrence coefficients. We also investigate the properties of the zeros of the orthogonal polynomials. Using Dyson’s Coulomb fluid approach together with the discrete system, we obtain the large $n$ asymptotic expansions of the recurrence coefficients, the sub-leading coefficient of the monic orthogonal polynomials, the Hankel determinant and the normalized constant for fixed $t>0$. It is found that all the asymptotic expansions are singular at $t=0$. We also study the long-time ($t\rightarrow+\infty$) asymptotics of these quantities explicitly for fixed $n\in\mathbb{N}$ from the Toda-type system.
@@ 1 Introduction
In the significant work [15], Chen and Its studied a singular linear statistic of the Laguerre unitary ensemble (LUE) inspired by an integrable quantum field theory at finite temperature [31]. This problem is equivalent to the study of the Hankel determinant and orthogonal polynomials for a singularly perturbed Laguerre weight
@@
The matrix model and Hankel determinant associated with the weight (1.1) have played an important role in the study of statistics for zeros of the Riemann zeta function [6] and the bosonic replica field theories [41]. Orthogonal polynomials and Hankel determinants for singularly perturbed Gaussian and Jacobi weights have also been extensively studied in the past few years; see, e.g., [7, 12, 33, 35, 39]. Matrix models with higher-order poles have been investigated in [1, 20, 21].
We would like to mention that orthogonal polynomials and Hankel determinants for the deformed measure $\mathrm{e}^{\frac{t}{x}}d\mu(x)$ have a very close relationship with the peakon solutions of the Camassa-Holm equation [3, 4, 5]. For more recent work in this respect, see [8, 9, 11]. Orthogonal polynomials and Hankel determinants for the singularly perturbed Laguerre weight (1.1) have been used to study the long-time asymptotics of the peakon solutions of the Camassa-Holm equation very recently by Chang, Eckhardt and Kostenko [10, Section 6].
@@
where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma\leq 2$ and deg $\tau$=1. For semi-classical orthogonal polynomials, they have a weight $w(x)$ that satisfies the Pearson equation (1.2), where $\sigma(x)$ and $\tau(x)$ are polynomials with deg $\sigma>2$ or deg $\tau\neq 1$. See, e.g., [45, Section 1.1.1]. It is not difficult to find that (1.1) is a semi-classical weight, since it satisfies the Pearson equation (1.2) with
@@
Please note that we use the notation $\alpha_{n},\beta_{n}$ for the recurrence coefficients (since this is standard), which should not cause confusion with the Laguerre weight parameter $\alpha$ as the latter is never subscripted with $n$.
From (1.3), (1.4) and (1.5), the recurrence coefficients also have the following expressions:
@@
Moreover, using the three-term recurrence relation (1.5) and with the aid of (1.7), we have the famous Christoffel-Darboux formula
$\sum_{j=0}^{n-1}\frac{P_{j}(x;t,\alpha)P_{j}(y;t,\alpha)}{h_{j}(t,\alpha)}=\frac{P_{n}(x;t,\alpha)P_{n-1}(y;t,\alpha)-P_{n}(y;t,\alpha)P_{n-1}(x;t,\alpha)}{h_{n-1}(t,\alpha)(x-y)},\qquad x\neq y.$
When $x=y$, the above formula becomes
$\sum_{j=0}^{n-1}\frac{P_{j}^{2}(x;t,\alpha)}{h_{j}(t,\alpha)}=\frac{P_{n}^{\prime}(x;t,\alpha)P_{n-1}(x;t,\alpha)-P_{n}(x;t,\alpha)P_{n-1}^{\prime}(x;t,\alpha)}{h_{n-1}(t,\alpha)},$
where denotes $\frac{d}{dx}$.
@@
where $K_{\nu}(z)$ is the modified Bessel function of the second kind of order $\nu$ and it has the integral representation [40, Equation (10.32.10)]
@@
The rest of the paper is organized as follows. In Section 2, by applying the ladder operators and compatibility conditions for the generalized Laguerre-type weight to our problem, we show that the orthogonal polynomials satisfy a second-order differential equation, and derive a discrete system and a Toda-type system for the recurrence coefficients. We also present the relations among the logarithmic derivative of the Hankel determinant, the sub-leading coefficient of the monic orthogonal polynomials, the recurrence coefficients and the auxiliary quantities. In Section 3, we obtain a mixed recurrence relation for the orthogonal polynomials and then study some properties of the zeros of the orthogonal polynomials. In Section 4, we investigate the large $n$ asymptotics of the recurrence coefficients $\alpha_{n}(t,\alpha)$ and $\beta_{n}(t,\alpha)$, the sub-leading coefficient $\mathrm{p}(n,t,\alpha)$, the Hankel determinant ${D_{n}}(t,\alpha)$ and the normalized constant ${h_{n}}(t,\alpha)$} for fixed $t>0$. We discuss the long-time ($t\rightarrow+\infty$) asymptotics of these quantities explicitly for fixed $n\in\mathbb{N}$ in Section 5.
@@ 2 Ladder operators, difference and differential equations
where $w_{0}(x)$ is continuously differentiable on $0,\infty)$ and all the moments of $w(x)$ exist. (Chen and Ismail’s ladder operator method requires $\alpha>0$.)
For our problem, the monic orthogonal polynomials $P_{n}(x)$ satisfy the ladder operator equations
@@
$B_{n}^{2}(x)+\mathrm{v}^{\prime}(x)B_{n}(x)+\sum_{j=0}^{n-1}A_{j}(x)=\beta_{n}A_{n}(x)A_{n-1}(x),$ ($S_{2}^{\prime}$)
where ($S_{2}^{\prime}$) is obtained from a suitable combination of ($S_{1}$) and ($S_{2}$). Next, we will use the ladder operators and associated compatibility conditions to analyze our problem.
@@
The expressions of ${A_{n}}(x)$ and ${B_{n}}(x)$ in the above lemma are coincident with [15, Lemma 2]. It can be seen that our derivation is more straightforward and the results are valid for $\alpha>-1,\;t\geq 0$.
@@
Remark 4.
We mention that Chen and Its [15] derived a system of difference equations for the auxiliary quantities $R_{n}$ and $r_{n}$ from (2.12)–(2.15):
$r_{n}+r_{n+1}=t-(2n+1+\alpha+R_{n})R_{n},$
$(r_{n}^{2}-tr_{n})(R_{n}+R_{n-1})=[nt-(2n+\alpha)r_{n}]R_{n}R_{n-1},$
which can be identified as the discrete Painlevé equations as pointed out by Van Assche [45, p. 58–59].
@@
The auxiliary quantity $R_{n}(t)$, related to the recurrence coefficient $\alpha_{n}(t)$ via (2.17), satisfies the Painlevé III  equation [32, (A.45.4 )]
@@
Then, the equation (2.35) is converted into the $\sigma$-form of the Painlevé III equation [29, (C.29)]
@@ 3 Zeros of orthogonal polynomials with the singularly perturbed Laguerre weight
$+\left(1-\frac{e_{n}}{\beta_{n-1}(t,\alpha)}\right)P_{n}(x;t,\alpha),$ (3.1)
@@
Following [19, Lemma 3.6], we have for our weight
@@
By using Christoffel’s formula (see, e.g., [27, Theorem 2.7.1]) and letting
@@
Let $x_{1,n}<x_{2,n}<\cdots<x_{n,n}$ denote the $n$ zeros of the orthogonal polynomials $P_{n}(x;t,\alpha)$ with respect to the singularly perturbed Laguerre weight (1.1). Then the zeros
@@
    $0<x_{1,n}<\alpha_{n-1}+\frac{d_{n}\beta_{n-1}}{e_{n}}<x_{n,n},$ (3.5)
@@
    are both increasing functions of $x$ on $\mathbb{R}^{+}$. Using Markov’s monotonicity theorem [43, Theorem 6.12.1], we have that the $j$th zero $x_{j,n}$ (for fixed $j=1,2,\ldots,n$) is an increasing function of both $t$ and $\alpha$.
@@
    which corresponds to $k=1$ and $G_{1}(x)=x-\alpha_{n-1}(t,\alpha)$ in [24, Equation (1)]. It follows from [24, Corollary 2.2] that $\alpha_{n-1}(t,\alpha)$ is an inner bound of the extreme zeros $x_{1,n}$ and $x_{n,n}$. The outer bounds $a_{n}$ and $b_{n}$ for the extreme zeros come from Ismail [27, Theorems 7.2.6 and 7.2.7] (see also Ismail and Li [28, Theorems 2 and 3]) using chain sequences.
@@
    Using the mixed recurrence relation (3.1) and following the similar analysis in [19, Theorem 3.9], it can be found that the zero of $G(x):=\frac{e_{n}}{\beta_{n-1}}\left(x-\alpha_{n-1}\right)-d_{n}$ (the coefficient of $P_{n-1}(x)$ in (3.1)) and the n2n-2 zeros of $P_{n-2}(x;t,\alpha+2)$ interlace with the $n$ zeros of $P_{n}(x;t,\alpha)$. It follows that the zero $\alpha_{n-1}+\frac{d_{n}\beta_{n-1}}{e_{n}}$ of $G(x)$ must lie between the two extreme zeros of $P_{n}(x;t,\alpha)$. Hence, we obtain another inner bound for the extreme zeros in (3.5).∎
@@
If $t=0$, our orthogonal polynomials $P_{n}(x;t,\alpha)$ are reduced to the classical (monic) Laguerre polynomials and $\alpha_{n}(0,\alpha)=2n+1+\alpha,\;\beta_{n}(0,\alpha)=n(n+\alpha)$. In this case, Theorem 3.2(iii) coincides with the result for classical Laguerre polynomials [27, Theorem 7.2.8] by choosing $c_{n}=4$ for $a_{n}$ and $c_{n}=4\cos^{2}\left(\frac{\pi}{n+1}\right)$ for $b_{n}$.
@@ 4 Large $n$ asymptotics
In the framework of Dyson’s Coulomb fluid approach [25], the eigenvalues of the Hermitian matrices from the unitary ensemble can be approximated by a continuous fluid with an equilibrium density when $n$ is sufficiently large. For our problem, it can be seen that the potential $\mathrm{v}(x)$ in (2.6) satisfies the condition that $x\mathrm{v}^{\prime}(x)$ is increasing on $\mathbb{R}^{+}$. This leads to the fact that the support of the equilibrium density is a single interval, denoted by (a,b),a>0(a,b),\;a>0; see [42, p. 199].
@@
Following the similar computations in [36, p. 406–407], multiplying by $\frac{1}{\sqrt{(b-x)(x-a)}}$ on both sides of (4.3) and integrating with respect to $x$ from $a$ to $b$, we have
@@
Our results for the large $n$ asymptotic expansions of the recurrence coefficients and the quantity $H_{n}(t)$ in Theorems 4.1 and 4.3 coincide with those in the work of Xu, Dai and Zhao [47, Corollary 1] using the Deift-Zhou steepest descent method for Riemann-Hilbert problems. Our method has the advantage of finding as many terms as one wants for the expansions. Furthermore, it can be seen that the large $n$ asymptotic expansions of the recurrence coefficients and the quantity $H_{n}(t)=t\frac{d}{{dt}}\ln{D_{n}}(t)$ are singular at the origin, which implies that the large $n$ asymptotic expansion of the Hankel determinant $D_{n}(t)$ can not be obtained directly by integrating the large $n$ asymptotic expansion of $H_{n}(t)/t$ from 0 to $t$. We derive the asymptotic expansion of the Hankel determinant in the following theorem. It was pointed out by Mezzadri and Mo [33, Section 1.1] that computing asymptotic formulae of Hankel determinants is a very important task—often a very difficult one—in many branches of mathematics and physics. The asymptotics of Hankel determinants depend crucially on the analytic properties of the weight $w(x)$. Usually, singular weights are the most challenging.
@@ 5 Long-time asymptotics
The long-time asymptotics (without higher-order terms) of the Hankel determinant in (5.7) has also been recently derived by Chang, Eckhardt and Kostenko [10, Lemma 6.7] employing different method, which has been used to obtain the long-time asymptotics of the peakon solutions of the Camassa-Holm equation. In addition, the constant (5.8) also appeared in the asymptotic expansion of the Hankel determinant for the generalized Airy weight as $t\to+\infty$ [38].